OpenStudy (anonymous):

solve for x... (1/x-3) + (1/x+3) = 10/x^2 - 9

OpenStudy (anonymous):

(1/x+3)+(1/x-3)=(10/x^2-9) [multiply by LCD (x-3)(x+3)] x-3+x+3=10 2x=10 /2 x=5 woot woot:DDD